OpenStudy (anonymous):

A dog owner is told by the veterinarian that his dog, Max, is overweight and needs to lose 4 to 5 pounds in order to live a healthy life. The dog weighs 126 pounds when he is put on a special diet to aid in the weight loss. The vet predicts that the diet will allow Max to lose 4 ounces per week. Max is brought to the vet every 2 weeks for a weight check. Write an equation to represent the doctor's prediction, and the relationship between Max's weight in pounds, y, and the number of times Max's weight is checked, x, in periods of two weeks . Reduce all answers to the nearest tenth of a number.

jigglypuff314 (jigglypuff314):

Hello and Welcome to OpenStudy! :) you would receive a quicker response by asking in the Mathematics Section instead http://www.openstudy.com/study#/groups/Mathematics @Jay_Blu

jigglypuff314 (jigglypuff314):

y = b + mx b = starting = 126 m = rate = decreasing at 0.5 pounds per two weeks 0.5 because... 16 ounces in 1 pound. 4 ounces per 1 week = 8 ounces per 2 weeks and 8 ounces = 0.5 pounds so y = b + mx plug in b = 126 and m = -0.5

OpenStudy (anonymous):

yay! thankyou

jigglypuff314 (jigglypuff314):

@Jay_Blu You're Welcome :)